OpenStudy (anonymous):

sum of 3^(1/n) from 1 to infinity

OpenStudy (anonymous):

really confused on how to do this would you use ratio test, comparisnt test?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

the series looks like this: \(3+3^{1/2}+3^{1/3}+\ldots\) an exponent of (1/n) implies n^th root and n^th root of "3" will always be greater than "1" and hence the sum will keep diverging. \[ L=\lim_{n\to\infty}3^{1/n}\\ \text{substitute}\quad u=1/n\qquad\therefore\text{as}\;n\to\infty,u\to0\\ L=\lim_{u\to0}3^u=1\ne0 \] by Limit test, the series diverges@@